# Hodge structures on conformal blocks

### v1

@@ 1.7. Organization of the paper
Section 3 is devoted to the definition of the notion of Hodge structure on a complex modular functor, to the statement of the main result (existence and uniqueness, 3.21) and to a discussion of the Frobenius algebra governing Hodge numbers and of the CohFT associated to characteristic classes of Hodge structures.
In Section 4, we use the results of the first paper [16] in this series and Theorem 3.21 to compute Hodge numbers of all $\mathrm{SU}(2)$ modular functors of level $2$ times an odd number. We also explain how one could use [7, 1.11] to compute Hodge numbers of modular functors associated to $\mathfrak{sl}_{n}$, and study gaps in Hodge decompositions (Section 4.3).
The relationship between a modular functor $\mathcal{V}$ and its associated ribbon category $\mathrm{C}_{\mathcal{V}}$ is the subject of Section 5. There, we prove that the functor $\mathcal{V}\mapsto\mathrm{C}_{\mathcal{V}}$ is fully-faithful.
@@ 2.1. Topological definition of a modular functor
2.2. Cobordism notation, genus 00, and rooted functor
@@
We know define rooted functors by allowing only connected genus 00 surfaces with exactly $1$ output, and gluings of an input of one surface to the output of another.
@@
A rooted functor is the data of functor:
@@
$\mathfrak{mod}^{0}\longrightarrow\mathfrak{root}$
from the category of genus 00 modular functors to the category of rooted functors.
@@ 2.3. Levels for modular functors
For modular functors, the following proposition is a direct consequence of the fact that, over a field, any finite-dimensional representation of the mapping class group of a surface of genus $g\geq 3$ sends Dehn twists to quasi-unipotent matrices [3, 2.5]. For genus 00 and rooted functors $\mathcal{V}$ with the assumption that the associated weakly ribbon category $\mathrm{C}_{\mathcal{V}}$ is ribbon, this is the Anderson-Moore-Vafa theorem ([12, 8.18.2]). See Sections 5.1 and 7.1 for the definitions of $\mathrm{C}_{\mathcal{V}}$ and ribbon category.
@@
Let $\mathcal{V}$ be a modular functor over an integral domain $R$ with set of colors $\Lambda$. Then the twists $t_{\mu}\in R^{\times}$ as in Remark 2.9 are roots of unity. The same conclusion holds for genus 00 and rooted functors $\mathcal{V}$ over integral domains when the associated weakly ribbon category $\mathrm{C}_{\mathcal{V}}$ is ribbon.
@@
Let $\mathcal{V}$ be a modular, genus 00 modular or rooted functor over a ring $R$ with set of colors $\Lambda$. A level for $\mathcal{V}$ is an integer $r$ such that $t_{\mu}^{r}=1$ for all $\mu\in\Lambda$. In particular, any modular functor over an integral domain admits a level. Similarly, genus 00 modular and rooted functors whose associated category is ribbon admit levels.
@@ 2.4. Geometric definition of a modular functor
However moduli spaces $\mathcal{M}_{g}^{n}$ are not compact/proper and the gluing maps are not algebraic (see [4, 6.2.11]). This makes the application of non-abelian Hodge theory on the $\mathcal{M}_{g}^{n}$ harder. Hence we replace them by proper DM stacks ¯$\overline{\mathcal{M}}_{g}^{n}(r)$, where $r$ is a level for the modular functor. With this approach, we get algebraic gluing maps, see Propositions 2.27 and 2.28.
@@
Notation 2.21.
Let ¯$\overline{\mathcal{M}}_{{g},{n}}(r)$ be the moduli space of connected stable nodal $r$-twisted curves of genus $g$ with $n$ distinct smooth marked points (see [10, 1.3]).
@@
Alternatively, ¯$\overline{\mathcal{M}}_{{g},{n}}(r)$ can be defined as follows. Let $D_{i}\subset\overline{\mathcal{M}}_{{g},{n}}$ for $i=1,\dotsc,k$ be the components of the boundary divisor. Then ¯$\overline{\mathcal{M}}_{{g},{n}}(r)$ is obtained from ¯$\overline{\mathcal{M}}_{{g},{n}}$ by taking $r$-th root stack independently on each $D_{i}$:
¯$\overline{\mathcal{M}}_{{g},{n}}(r)=\overline{\mathcal{M}}_{{g},{n}}[D_{1}/r]\times_{\overline{\mathcal{M}}_{{g},{n}}}\dotsb\times_{\overline{\mathcal{M}}_{{g},{n}}}\overline{\mathcal{M}}_{{g},{n}}[D_{k}/r].$
With this description, it is clear that the fundamental groupoid of ¯$\overline{\mathcal{M}}_{{g},{n}}(r)$ is equivalent to the quotient $\mathrm{PMod}(S_{g,n})/\langle T_{\delta}^{r}\mid\delta\rangle$ of the pure mapping class group of the $n$ times punctured genus $g$ surface by all $r$-th powers of Dehn twists.
@@
Again, as in the topological case, one can give similar geometric definitions of genus 00 modular functor of level $r$ and rooted functor of level $r$: for the genus 00 case we restrict the above definition to $g=0$ and remove the (G-sep) axiom; for rooted functors, we replace the gluing axioms by the gluing axiom (G-nonsep) in cobordism notation and in the permutation axiom allow only permutations of inputs. We now define isomorphisms of geometric modular functors.
@@
Again, we have a similar result for genus 00 modular functors of level $r$ and rooted functors of level $r$.
@@ 2.5. Anomalous modular functors
One defines isomorphisms of topological and geometric modular functors as in Definitions 2.10 and 2.30. Then again, by [4, 6.7.12], one has the following.
@@ 3.1.1. Hodge structures on flat bundles
3.1.2. Semisimplicity of quantum representations
Any local system that supports a CVHS is semisimple ([23, 4]). Semisimplicity will be an assumption in our main result (3.21). Here we explain why this will be satisfied by most examples (such as conformal blocks coming from Lie algebras).
Definition 3.8.
Let $\mathcal{V}$ be a modular functor. It will be said to have semisimple quantum representations if for every $g$, $n$ and $\underline{\lambda}\in\Lambda^{n}$, the local system $\mathcal{V}_{g}(\lambda_{1},\dotsc,\lambda_{n})$ is semisimple. Similarly for genus 00 modular functors and rooted functors.
@@ 3.1.2. Semisimplicity of quantum representations
The author does not know of any modular functor over $\mathbb{C}$ that does not satisfy semisimplicity of quantum representations.
The following proposition is a direct consequence of Galois invariance of semisimplicity and provides a large family of modular functors with semisimple quantum representations.
Proposition 3.10.
Let $K$ be a number field and $\mathcal{V}$ a modular, genus 00 modular or rooted functor defined over $K$. Assume that there exists an embedding $\iota_{0}:K\hookrightarrow\mathbb{C}$ such that the bundles of the functor $\mathcal{V}^{\iota_{0}}$ over $\mathbb{C}$ obtained by base change are all unitary. Then $\mathcal{V}$ has semisimple quantum representations. In particular, for any embedding $\iota:K\hookrightarrow\mathbb{C}$, $\mathcal{V}^{\iota}$ has semisimple quantum representations.
The following theorem shows that any unitary complex modular functor whose associated category is ribbon gives rise to some Galois conjugates, that may then not be unitary and thus be more interesting from the point of view of Hodge theory.
Theorem 3.11 (7.14).
Let $\mathcal{V}$ be a complex modular, genus 00 modular or rooted functor whose associated weakly ribbon category $\mathrm{C}_{\mathcal{V}}$ is ribbon (see Sections 5.1 and 7.1). Then there exists a number field $K\subset\mathbb{C}$ such that $\mathcal{V}$ is definable over $K$.
Remark 3.12.
Complex modular functors coming from complex finite dimensional simple Lie algebras $\mathfrak{g}$ by the conformal blocks construction are known to be unitary (see [6]). Hence the discussion above applies. In the case $\mathfrak{g}=\mathfrak{sl}_{2}$, up to a change of sign in the twists of the ribbon structure, the associated modular functors are equivalent to the $\mathrm{SU}(2)$ modular functors. See Sections 4 and 4.2 below for more details.
@@ 3.1.3. Hodge structures on modular functors
One can give similar definitions for genus 00 modular and rooted functors.
@@
Let $\mathcal{V}$ be a complex modular, genus 00 modular or rooted functor whose associated weakly ribbon category $\mathrm{C}_{\mathcal{V}}$ is ribbon (see Sections 5.1 and 7.1). Then any weight function on $\mathcal{V}$ is exact.
@@
Similarly, one defines the notions of complex Hodge structure on a genus 00 modular functors and on a rooted functor, where, for rooted functors, we omit the assumption on $\mathcal{V}_{0}(0,0)$ as it is not defined in that case.
@@
Let $\mathcal{V}$ be a complex modular, genus 00 modular or rooted functor, with set of colors $\Lambda$. Then $\mathrm{Shifts}(\Lambda)$ acts on the set of complex Hodge structures on $\mathcal{V}$, and the action of its subspace $\mathrm{Shifts}^{0}(\Lambda)$ preserves weight functions.
@@ 3.2. Existence and uniqueness
Let $\mathcal{V}$ be a complex modular, genus 00 modular or rooted functor, with semisimple quantum representations and such that the associated weakly ribbon category $\mathrm{C}_{\mathcal{V}}$ is ribbon. Denote by $\Lambda$ its set of colors. Then for any weight function $w$, $\mathcal{V}$ admits a complex Hodge structure whose weights are given by $w$. Moreover, $\mathrm{Shifts}^{0}(\Lambda)$ acts transitively on the set of complex Hodge structures on $\mathcal{V}$ with weight function $w$.
For a definition of $\mathrm{C}_{\mathcal{V}}$ and ribbon categories, see Sections 5.1 and 7.1.
@@
Let $\mathcal{V}$ be a complex modular functor with semisimple quantum representations such that $\mathrm{C}_{\mathcal{V}}$ is ribbon. Then any complex Hodge structure on the rooted functor 666which is a genus 00 structure! associated to $\mathcal{V}$ extends uniquely to the whole modular functor. In particular, if the Hodge numbers are known for the associated rooted functor, then they are known for the modular functor.
@@ 3.3. The Frobenius algebra governing the Hodge numbers
Let $\mathcal{V}$ be a rooted functor over $\mathbb{C}$ with a Hodge structure. Denote by $\Lambda$ its set of colors. Define $V$ as the free $\mathbb{Z}[u,v]_{\mathrm{rat}}$-module with basis $\Lambda$. Let $\star$ be the symmetric bilinear operation on $V$ defined by:
@@
Similarly, for $\mathcal{V}$ a rooted functor with Hodge structure, the Hodge polynomial $e(\mathcal{V}(\mu;\lambda_{1},\dotsc,\lambda_{n}))$ is equal to the coefficient of $[\mu]$ in $[\lambda_{1}]\star\dotsb\star[\lambda_{n}]$.
The idea that Hodge numbers of rooted functors could be organized in an algebra appeared independently in the recent paper [7] of Belkale, Fakhruddin and Mukhopadhyay. See Sections 1.6 and 4.2 for more comments on their paper.

### v2

@@ 1.7. Organization of the paper
Section 3 is devoted to the definition of the notion of Hodge structure on a complex modular functor, to the statement of the main result (existence and uniqueness, 3.19) and to a discussion of the Frobenius algebra governing Hodge numbers and of the CohFT associated to characteristic classes of Hodge structures.
In Section 4, we use the results of the first paper [18] in this series and Theorem 3.19 to compute Hodge numbers of all $\mathrm{SU}(2)$ modular functors of level $2$ times an odd number. We also explain how one could use [6, 1.11] to compute Hodge numbers of modular functors associated to $\mathfrak{sl}_{n}$, and study gaps in Hodge decompositions (Section 4.3).
The relationship between a modular functor $\mathcal{V}$ and its associated ribbon fusion category $\mathrm{C}_{\mathcal{V}}$ is the subject of Section 5. There, we prove that the functor $\mathcal{V}\mapsto\mathrm{C}_{\mathcal{V}}$ is fully-faithful.
@@ 2.1. Topological definition of a modular functor
2.2. Cobordism notation, genus 00, ribbon and braided functors
@@
We know define ribbon functors by allowing only connected genus 00 surfaces with exactly $1$ output, and gluings of an input of one surface to the output of another.
@@
A ribbon functor is the data of functor:
@@
Definition 2.15.
For $\Lambda$ a finite set with preferred element 00, let $\mathrm{Disk}_{\Lambda}^{b}$ be the fundamental groupoid of the space [^4] The connected components of this space are parametrized by the number of disks embedded and the colorings. Each component is parametrized by the possible centers $x_{i}$ and radii $r_{i}$. of finite configurations of disjoint $\Lambda$-colored open disks $\{(\mathbb{D}(x_{i},r_{i}),\lambda_{i})\mid i\}$ in a $\Lambda$-colored unit disk $(\overline{\mathbb{D}}(0,1),\mu)$. We may see an object in $\mathrm{Disk}_{\Lambda}^{b}$ as an object of $\mathrm{Disk}_{\Lambda}$ by considering $\overline{\mathbb{D}}(0,1)\setminus\bigsqcup_{i}\mathbb{D}(x_{i},r_{i})$. By replacing $\mathrm{Disk}_{\Lambda}$ by $\mathrm{Disk}_{\Lambda}^{b}$ in the Definition 2.14, one gets the definition of braided functor.
The essential difference between $\mathrm{Disk}_{\Lambda}^{b}$ and $\mathrm{Disk}_{\Lambda}$ is that Dehn twists around input boundary components are morphisms in $\mathrm{Disk}_{\Lambda}$ but not in $\mathrm{Disk}_{\Lambda}^{b}$. Hence, for $\mathcal{V}$ a braided functor, the group $\mathrm{Aut}((D^{n},\mu;\underline{\lambda}))$ acting on $\mathcal{V}(D^{n},\mu;\underline{\lambda})$ is a braid group while, for $\mathcal{V}$ a ribbon functor, it is a (larger) framed braid group.
@@
$\mathfrak{mod}^{0}\longrightarrow\mathfrak{ribfun}$
from the category of genus 00 modular functors to the category of ribbon functors. Moreover, the functor $\mathrm{Disk}_{\Lambda}^{b}\rightarrow\mathrm{Disk}_{\Lambda}$ induces a functor:
$\mathfrak{ribfun}\longrightarrow\mathfrak{braidfun}$
from the category of ribbon functors to the category of braided functors.
@@ 2.3. Levels for modular functors
For modular functors, the following proposition is a direct consequence of the fact that, over a field, any finite-dimensional representation of the mapping class group of a surface of genus $g\geq 3$ sends Dehn twists to quasi-unipotent matrices [3, 2.5]. For genus 00 modular functors and ribbon functors $\mathcal{V}$ this is the Anderson-Moore-Vafa theorem ([12, 8.18.2]). The case of braided functors requires an extra argument, that was communicated to the author by Pavel Etingof, and that we now sketch. A braided fusion category $\mathcal{C}$ admits a pivotalization $\widetilde{\mathcal{C}}\rightarrow\mathcal{C}$. It is a spherical category (see [12, 7.21.10]) whose simple objects map $2$ to $1$ onto those of $\mathcal{C}$. Moreover, the braiding on $\mathcal{C}$ lifts to $\widetilde{\mathcal{C}}$. As $\widetilde{\mathcal{C}}$ is spherical, braided and fusion, it is ribbon (see [21, A.4]). In particular the braiding on $\widetilde{\mathcal{C}}$ has finite order on each pair of objects. Hence the braiding on $\mathcal{C}$ has the same property. See Sections 5.1 and 7.1 for the definitions of braided and ribbon fusion categories, and their relation to braided and ribbon functors.
@@
Let $\mathcal{V}$ be a modular functor over an integral domain $R$ with set of colors $\Lambda$. Then the twists $t_{\mu}\in R^{\times}$ as in Remark 2.9 are roots of unity. The same conclusion holds for genus 00 modular functors and ribbon functors over integral domains.
Let $\mathcal{V}$ a braided functor over an integral domain $R$ with set of colors $\Lambda$. Then with the notation $\beta_{\lambda;\mu,\nu}:\mathcal{V}(\lambda;\mu,\nu)\rightarrow\mathcal{V}(\lambda;\nu,\mu)$ for the braidings, the eigenvalues of the $\beta_{\lambda;\mu,\nu}\circ\beta_{\lambda;\mu,\nu}$ are roots of unity.
@@
Let $\mathcal{V}$ be a modular, genus 00 modular or ribbon functor over a ring $R$ with set of colors $\Lambda$. A level for $\mathcal{V}$ is an integer $r$ such that $t_{\mu}^{r}=1$ for all $\mu\in\Lambda$. For $\mathcal{V}$ a braided functor over a ring $R$ with set of colors $\Lambda$, a level is an integer $r$ such that each $(\beta_{\lambda;\mu,\nu}\circ\beta_{\lambda;\mu,\nu})^{r}$ is the identity.
By Proposition 2.17, all modular, genus 00 modular, ribbon or braided functors over integral domains admit levels.
@@
One can provide similar definitions of $\mathrm{Surf}_{\Lambda}^{0}(r)$, $\mathrm{Disk}_{\Lambda}(r)$ and $\mathrm{Disk}_{\Lambda}^{b}(r)$, for which analogs of Proposition 2.20 hold.
@@ 2.4. Geometric definition of a modular functor
However moduli spaces $\mathcal{M}_{g}^{n}$ are not compact/proper and the gluing maps are not algebraic (see [4, 6.2.11]). This makes the application of non-Abelian Hodge theory on the $\mathcal{M}_{g}^{n}$ harder. Hence we replace them by proper DM stacks ¯$\overline{\mathcal{M}}_{g}^{n}(r)$, where $r$ is a level for the modular functor. With this approach, we get algebraic gluing maps, see Propositions 2.28 and 2.29.
@@
Notation 2.22.
Let ¯$\overline{\mathcal{M}}_{{g},{n}}(r)$ be the moduli space of connected stable nodal $r$-twisted curves of genus $g$ with $n$ distinct smooth marked points (see [9, 1.3]).
@@
Alternatively, ¯$\overline{\mathcal{M}}_{{g},{n}}(r)$ can be defined as follows. Let $D_{i}\subset\overline{\mathcal{M}}_{{g},{n}}$ for $i=1,\dotsc,k$ be the components of the boundary divisor. Then ¯$\overline{\mathcal{M}}_{{g},{n}}(r)$ is obtained from ¯$\overline{\mathcal{M}}_{{g},{n}}$ by taking $r$-th root stack independently locally on each $D_{i}$:
¯$\overline{\mathcal{M}}_{{g},{n}}(r)=\overline{\mathcal{M}}_{{g},{n}}\left[\sum_{i}\frac{D_{i}}{r}\right].$
See [9, 2.3, 4.5] for more details on what is meant by independent root stack on a normal crossing divisor with self-crossings in this context. With this description, it is clear that the fundamental groupoid of ¯$\overline{\mathcal{M}}_{{g},{n}}(r)$ is equivalent to the quotient $\mathrm{PMod}(S_{g,n})/\langle T_{\delta}^{r}\mid\delta\rangle$ of the pure mapping class group of the $n$ times punctured genus $g$ surface by all $r$-th powers of Dehn twists.
@@
Again, as in the topological case, one can give similar geometric definitions of genus 00 modular functor of level $r$ and ribbon functor of level $r$: for the genus 00 case we restrict the above definition to $g=0$ and remove the (G-sep) axiom; for ribbon functors, we replace the gluing axioms by the gluing axiom (G-nonsep) in cobordism notation and in the permutation axiom allow only permutations of inputs.
As for braided functors, one needs to work with slight variations of the moduli spaces.
Notation 2.31.
For $n\geq 1$, denote by $^{1}\overline{\mathcal{M}}_{0,{n}}(r)$ the moduli space of connected stable nodal $r$-twisted curves of genus 00 with $n$ distinct smooth points and $1$ order $r$ stacky point. and a section at the stacky point. We will use the convention $^{1}\overline{\mathcal{M}}_{0,{1}}(r)=*$.
Remark 2.32.
Let us mark the points in ¯$\overline{\mathcal{M}}_{{0},{n+1}}(r)$ from 00 to $n$. The space $^{1}\overline{\mathcal{M}}_{0,{n}}(r)$ can de identified with $\Sigma_{0}\subset\overline{\mathcal{C}}_{{0},{n+1}}(r)$, the 00-th stacky marked point $\Sigma_{0}$ in the universal curve $\overline{\mathcal{C}}_{{g},{n}}(r)\rightarrow\overline{\mathcal{M}}_{{g},{n}}(r)$. Its fundamental group can be identified with $\mathrm{PMod}(D_{n})/\langle T_{\delta}^{r}\mid\delta\rangle$, where $D_{n}$ is the closed disk with $n$ points removed. The pure mapping class group $\mathrm{PMod}(D_{n})$ is the pure braid group $\mathrm{PB}_{n}$.
Replacing the ¯$\overline{\mathcal{M}}_{0}^{n+1}(r)$ with the $^{1}\overline{\mathcal{M}}_{0,{n}}(r)$ in the definition of geometric ribbon functor, one gets the definition of geometric braided functor. For details, the reader can find a full definition in [19, 2.20].
We now define isomorphisms of geometric modular functors.
@@
Again, we have similar results for genus 00 modular functors of level $r$, ribbon functors of level $r$ and braided functors of level $r$.
@@ 2.5. Anomalous modular functors
One defines isomorphisms of topological and geometric modular functors as in Definitions 2.10 and 2.33. Then again, by [4, 6.7.12], one has the following.
@@ 3.1.1. Hodge structures on flat bundles
3.1.2. Semisimplicity and definition over number fields
As mentioned in Remark 1.1, the main result (Theorem 3.19) relies on knowing that conformal blocks are semisimple. The semisimplicity of conformal blocks in general is proved in the separate paper [20].
Theorem 3.8 ([20, 1.1]).
Let $\mathcal{V}$ be a geometric modular functor. Then, for any $g,n\geq 0$, $(g,n)\neq(0,0),(0,1),(1,0)$, and $\underline{\lambda}\in\Lambda$, the conformal block $(\mathcal{V}_{g}(\underline{\lambda}),\nabla)$ is semisimple. The same holds for genus 00 modular, ribbon or braided functors.
The following is a known consequence of Ocneanu rigidity.
Theorem 3.9 (7.14).
Let $\mathcal{V}$ be a topological complex modular, genus 00 modular, ribbon or braided functor. Then there exists a number field $K\subset\mathbb{C}$ such that $\mathcal{V}$ is definable over $K$.
The theorem above shows that any unitary complex modular functor gives rise to some Galois conjugates, that may then not be unitary and thus be more interesting from the point of view of Hodge theory. This is what happens in the case of complex modular functors coming from complex finite dimensional simple Lie algebras $\mathfrak{g}$. They are known to be unitary (see [5]), and hence carry trivial uninteresting Hodge structures. However, in the topological definition, they are defined over cyclotomic fields and their Galois conjugates have non-trivial Hodge structures (see Sections 1.5 and 4).
@@ 3.1.2. Semisimplicity and definition over number fields
In the case $\mathfrak{g}=\mathfrak{sl}_{2}$, up to a change of sign in the twists of the ribbon structure, the associated modular functors are equivalent to the $\mathrm{SU}(2)$ modular functors. See Sections 4 and 4.2 below for more details.
@@ 3.1.3. Hodge structures on modular functors
One can give similar definitions for genus 00 modular and ribbon functors.
@@
Let $\mathcal{V}$ be a complex modular, genus 00 modular, ribbon or braided functor. Then any weight function on $\mathcal{V}$ is exact.
@@
Similarly, one defines the notions of complex Hodge structure on a genus 00 modular functors and on a ribbon or braided functor, where, for ribbon and braided functors, we omit the assumption on $\mathcal{V}_{0}(0,0)$ as it is not defined in that case.
@@
Let $\mathcal{V}$ be a complex modular, genus 00 modular or ribbon functor, with set of colors $\Lambda$. Then $\mathrm{Shifts}(\Lambda)$ acts on the set of complex Hodge structures on $\mathcal{V}$, and the action of its subspace $\mathrm{Shifts}^{0}(\Lambda)$ preserves weight functions.
@@ 3.2. Existence and uniqueness
Let $\mathcal{V}$ be a complex modular, genus 00 modular, ribbon or braided functor. Denote by $\Lambda$ its set of colors. Then for any weight function $w$, $\mathcal{V}$ admits a complex Hodge structure whose weights are given by $w$. Moreover, $\mathrm{Shifts}(\Lambda)$ acts transitively on the set of complex Hodge structures on $\mathcal{V}$.
@@
Let $\mathcal{V}$ be a complex modular functor. Then any complex Hodge structure on the braided functor 666which is a genus 00 structure! associated to $\mathcal{V}$ extends uniquely to the whole modular functor. In particular, if the Hodge numbers are known for the associated braided functor, then they are known for the modular functor.
@@ 3.3. The Frobenius algebra governing the Hodge numbers
Let $\mathcal{V}$ be a ribbon functor over $\mathbb{C}$ with a Hodge structure. Denote by $\Lambda$ its set of colors. Define $V$ as the free $\mathbb{Z}[u,v]_{\mathrm{rat}}$-module with basis $\Lambda$. Let $\star$ be the symmetric bilinear operation on $V$ defined by:
@@
Similarly, for $\mathcal{V}$ a ribbon or braided functor with Hodge structure, the Hodge polynomial $e(\mathcal{V}(\mu;\lambda_{1},\dotsc,\lambda_{n}))$ is equal to the coefficient of $[\mu]$ in $[\lambda_{1}]\star\dotsb\star[\lambda_{n}]$.
The idea that Hodge numbers of braided functors could be organized in an algebra appeared independently in the recent paper [6] of Belkale, Fakhruddin and Mukhopadhyay. See Sections 1.6 and 4.2 for more comments on their paper.